# Competitive Equilibrium for Electricity Markets with Spatially Flexible Loads

### v2

@@ VI Case Study
We study an integrated system comprising the power grid, transportation network, and datacenter network. This case study illustrates how the framework developed in Section II can be applied to real-world settings, and examines how the resulting GCE captures the interactions and mutual influences among these networks and affects welfare outcomes.
@@ VI-A Configurations of the Power and Multi-FL Systems
The power system configuration is modified from the setting in [27], built upon an 11-zone aggregation of the NYISO network that includes generation data, network parameters, baseline load profiles, and inter-zone distance information. The transportation and datacenter networks are electrically connected to the same power system. The transportation network is adapted from the Sioux Falls road network [35], consisting of 24 nodes, 76 edges. Each transportation node is equipped with an EV charging station that connects electrically to one of the 11 zones in the NYISO-based power system, with each zone containing at least one charging stations; these connections determine $\mathbf{A}^{\mathrm{CB}}$. Five datacenters are placed according to [27] and connected to their zonal power-system buses, deciding $\mathbf{A}^{\mathrm{DB}}$. The GCE is evaluated over discrete one-hour market-clearing periods, with all energy-use activities within each interval assumed to occur simultaneously. The configuration details can be found in Appendix 7.
@@ VI-A1 Electrified transportation system
The transportation system is modified from [35], comprising 182 OD pairs with OD demand $\bm{\rho}$ and hosting a total of 158,080 EVs. For each OD pair, the EV driver chooses between two candidate routes with the shortest free-flow travel times: one assuming charging at the origin and the other assuming charging at the destination, which determines $\mathbf{A}^{\mathrm{CR}}$. The travel cost function on edge $e$ is given by $c_{e}(y_{e}(\mathbf{x}))=c_{e}^{0}\big(1+\kappa_{e}y_{e}(\mathbf{x})\big)$, where $c_{e}^{0}$ represents the base cost, computed as the free-flow travel time multiplied by a value of time of \$17.09/hour [36]. The coefficient $\kappa_{e}$ reflects the marginal effect of traffic volume and is obtained by dividing the delay parameter by the edge capacity in [35]. Each EV requires an identical charging amount $q^{\mathrm{EV}}\in[1,20]$ kWh, corresponding to the energy delivered during one hour of Level 1-2 charging [37].
@@
VI-A2 Datacenter network
The system includes 11 AI companies, each associated with one of the 11 zones in New York. Each company flexibly allocates its computational workload across the five datacenters, with its utility function following the quadratic form (23). The specifications of parameters in the utility function are documented in Appendix 7. Each standardized workload (e.g., an AI inference/training batch) has an energy requirement of $q^{\mathrm{DC}}\in[0,10]$ kWh. Each AI company optimizes the allocation of its fixed total workload across datacenters to maximize its net payoff, and the total amount of workloads is 364,803.
@@ VI-B GCE-Based Interaction and Welfare Analysis
Using the GCE and SWM formulations developed in Section III, we compute the FL desions, generator dispatch, and LMPs $(\mathbf{g}^{\star},\mathbf{s}^{\star},\bm{\lambda}^{\star})$ under GCE and compare them with the corresponding social welfare optimal outcomes. We then vary key parameters, namely, the transmission line capacities, the EV charging amount $q^{\mathrm{EV}}$ and the energy demand per workload $q^{\mathrm{DC}}$, to examine their impacts on the performance of each subsystem. For comparison, we consider two experimental settings: the congested case, based on the original dataset where limited line capacity causes congestion even without interconnection of the FL systems, and the uncongested case, a synthetic scenario in which all line-capacity constraints are relaxed to ensure an uncongested power system.
@@
We examine the impact of spatial flexibility of the datacenter network by varying the number of available datacenters from 1 to 5 in Fig. 6 with the total workload from the AI companies fixed, where fewer datacenters indicate less flexibility for AI companies to allocate their workloads. The one- and two-datacenter cases are infeasible because both sites lie within the power system’s most congested zone. In Fig. 6(a), increasing the number of datacenters decreases generation cost, as greater spatial flexibility enables workload redistribution toward more efficient and less congested regions of the grid. Spatial flexibility of the datacenter network likewise reduces EVs’ total travel cost under both GCE and SWM frameworks in Fig. 6(b) by flattening LMP disparities, thus mitigating traffic congestion in the transportation network induced by drivers chasing cheaper charging prices. As shown in Fig. 6(c), computing utility improves with spatial flexibility due to reduced queuing delays. When datacenter locations are restricted, workloads concentrate in a few sites, resulting in negative utility from excessive waiting. Overall, Fig. 6 demonstrates that expanding spatial flexibility benefits all coupled infrastructures even in a self-interested GCE model.
@@ VII Conclusion
Appendix A Proofs in Section III
@@
Since $J^{\mathrm{SW}}$ and $C$ are convex and $\mathcal{F}$ is closed and convex, the standard subgradient optimality condition for minimizing a convex function over a convex set (see [38]) implies $(\mathbf{s}^{\star},\mathbf{g}^{\star})$ is optimal for $\min_{(\mathbf{s},\mathbf{g})\in\mathcal{F}}J^{\mathrm{SW}}(\mathbf{s})+C(\mathbf{g})$ if and only if there exist $\bm{\mu}\in\partial J^{\mathrm{SW}}(\mathbf{s}^{\star}),\ \bm{\nu}\in\partial C(\mathbf{g}^{\star})$ such that
@@
Appendix B Proofs in Section IV
@@
which is exactly the UE condition (14). Conversely, any $\mathbf{x}^{\star}$ satisfying (14) admits such multipliers $\mu_{k},\nu_{r}$ and therefore satisfies the KKT conditions, proving necessity and sufficiency. This establishes the equivalence between the optimization problem (15) and the UE condition (14). A similar derivation appears in the classical transportation equilibrium theory [32], where the charging component is absent. ∎
@@
TABLE I: Illustrative Example: Non-congested Case
@@
    No congestion $\bigl(f^{\max}\geq g_{2}^{\max}-(a-b)q/a\bigr)$. The line is not congested in both programs. Table I demonstrates the results under this parameter regime. We have $s_{1}^{\mathrm{GCE}}=2s_{1}^{\mathrm{SWM}}$ while the LMPs remain uniform, $\lambda_{1}^{\mathrm{GCE}}=\lambda_{2}^{\mathrm{GCE}}=\lambda_{1}^{\mathrm{SWM}}=\lambda_{2}^{\mathrm{SWM}}$. This result is an example for Lemma 3-i), where $(\mathrm{P}_{\mathrm{GCE}})$ does not support social welfare.
@@
Appendix C Proofs in Section V
@@
Since each utility $u_{k}$ is differentiable and concave and each feasible set $\mathcal{X}^{(k)}$ is convex and compact, the players’ optimization problem in (20) is convex. Its KKT conditions are therefore necessary and sufficient. Collecting the $K$ stationarity conditions yields exactly the VI in (21); Conversely, any $\mathbf{x}^{\star}$ solving (21) satisfies every player’s KKT system and is thus a Nash equilibrium. This one-to-one correspondence follows from Proposition 1.3.4 of [39].
Now suppose the cross-partial symmetry condition (22) holds. By standard results in potential game theory [33], there exists a continuously differentiable scalar function $\Phi(\mathbf{x})$ satisfying $\nabla_{\mathbf{x}^{(k)}}\Phi=-\nabla_{\mathbf{x}^{(k)}}u_{k}$ for all $k$. Substituting this relation into (21) gives
@@
Appendix D Non-Potential Game Case for Lemma 4
Theorem 3 (GCE for an Non-Potential Game).
[…]
Proof.
Fix an LMP vector $\bm{\lambda}\in\mathbb{R}^{N}$ and define the augmented operator $\mathbf{F}_{\bm{\lambda}}(\mathbf{x})\;:=\;\mathbf{F}(\mathbf{x})\;+\;\bigl(\mathbf{A}^{\mathrm{FL}}\bigr)^{\!\top}\bm{\lambda}.$ The VI problem (21) can be denoted as $\operatorname{VI}(\mathcal{X},\mathbf{F}_{\bm{\lambda}})$. Since $\mathbf{F}$ is strictly monotone on the compact, convex set $\mathcal{X}$, $\mathbf{F}_{\bm{\lambda}}$ is strictly monotone as well. Therefore, $\operatorname{VI}(\mathcal{X},\mathbf{F}_{\bm{\lambda}})$ has a unique solution according to Theorem 2.3.3 in [39]. The merit function for $\operatorname{VI}(\mathcal{X},\mathbf{F}_{\bm{\lambda}})$ is as follows:
@@ Appendix D Non-Potential Game Case for Lemma 4
Next, we illustrate the relationship between the characterization of GCE and the solution of $\operatorname{VI}(\Omega,\mathbf{F}_{\mathbf{C}})$. Since $C$ is convex and $\mathbf{F}$ is monotone, $\mathbf{F}_{\mathbf{C}}$ is a monotone operator on the convex set $\Omega$. By Proposition 1.3.4 of [39], a pair $(\mathbf{x}^{\star},\mathbf{g}^{\star})$ solves $\operatorname{VI}(\Omega,\mathbf{F}_{\mathbf{C}})$ if and only if there exists a multiplier $\bm{\lambda}^{\star}$ associated with the coupling constraints $\mathbf{g}-\bm{\ell}-\mathbf{A}^{\mathrm{FL}}\mathbf{x}\!\in\!\mathcal{P}$ such that the triplet $(\mathbf{x}^{\star},\mathbf{g}^{\star},\bm{\lambda}^{\star})$ satisfies the KKT conditions for the VI, which consist of three parts:
@@
Appendix E Case Study Configuration Details
The 11 zones of the power system, indexed by A-K, are shown in Fig. 7(a). Each zone represents a major operating region (e.g., WEST, CENTRL, N.Y.C), and the transmission lines connecting them indicate the network topology of the regional grid. The coefficient of the quadratic term in the generation cost is set to be ten times larger than that in [27], reflecting a steeper marginal generation cost profile.
The transportation system, shown in Fig. 7(b), consists of 24 nodes representing charging stations as well as the origins and destinations of EV trips. These nodes are grouped into the same 11 regions (A-K) as the power system, with shaded blue areas denoting their geographic correspondence and coupling to the power zones. Arrowed lines indicate the bidirectional links between nodes, representing the road connections that form the transportation network. To better capture real-world congestion effects, the travel demand of each OD pair is adjusted, the free-flow travel time is tripled, and the delay parameter $\kappa_{e}$ is amplified by a factor of ten. The adjustment of the generator quadratic cost coefficients and delay parameters is introduced to increase the sensitivity of marginal generation cost/per driver travel cost to the relocation of traffic flow.
In Fig. 7(c), the geographic locations of the 11 AI companies and 5 datacenters are marked on the map. The solid lines link each AI company and datacenter to their respective regional indices (A-K), indicating their regional associations. Specifically, five datacenters are located in the WEST, CENTRL, NORTH, MILLWD, and N.Y.C regions, consuming electricity locally within their respective zones. For the datacenter network, the dashed lines shown for one representative company illustrate that each AI company can connect to and dispatch workloads across all five datacenters, with different colors corresponding to different datacenters.
TABLE II: Parameters in AI Companies’ Utility Functions (23).
Param Description Value
$d_{j}^{(k)}$ Distance from company $k$ to datacenter $j$ based on [27]
$\zeta_{k}$ Total workload of user $k$ $10^{4}\cdot\mathcal{U}(1,5)$
In (23), each component $\eta_{k,j}$ of the baseline utility vector represents company $k$’s willingness to pay (or preference) for datacenter $j$, denoted by $\tau^{(k)}_{j}$, minus a communication cost proportional to the distance $\gamma_{k}d^{(k)}_{j}$ between them. The quadratic cost terms arise from datacenter queuing effects [34]. At each datacenter $j$, a per-job waiting cost increases proportionally with the total load, $\sum_{\tilde{k}}x_{j}^{(\tilde{k})}$, governed by a queuing coefficient $\alpha_{j}\geq 0$. The total waiting cost for company $k$ is then given by $\sum_{j}\alpha_{j}x_{j}^{(k)}\sum_{\tilde{k}}x_{j}^{(\tilde{k})}$, establishing the interaction between the workloads of competing companies. This interaction directly defines the off-diagonal blocks as $\mathbf{M}_{k,\tilde{k}}=\operatorname{diag}(\alpha_{1},\ldots,\alpha_{N^{\mathrm{D}}})$ for every $\tilde{k}\neq k$, and the row block matrix $\mathbf{M}_{k,-k}$ horizontally concatenates all the individual interaction matrices $\mathbf{M}_{k,\tilde{k}}$. This queuing cost is then combined with an internal cost that captures diminishing marginal returns, modeled with parameter $\omega_{k}>0$ in the quadratic form $\frac{1}{2}\omega_{k}\sum_{j}(x_{j}^{(k)})^{2}$. Then the diagonal blocks satisfy $\mathbf{M}_{k,k}=\operatorname{diag}(2\alpha_{1}+\omega_{k},\ldots,2\alpha_{N^{\mathrm{D}}}+\omega_{k})$. This formulation guarantees the utility function’s concavity, as the requirement $\mathbf{M}_{k,k}\succeq 0$ is always satisfied. The parameters of the utility functions in our simulations are selected according to the aforementioned criteria and summarized in Table II, where $\mathcal{U}(\cdot)$ denotes the uniform distribution. Once the random components of the parameters in the table are determined, these parameters remain fixed throughout all experiments for consistency.
When we vary the number of datacenters from 1 to 5 for assessing the impact of spatial flexibility, we connect each AI company to an expanding set of datacenters, starting from the datacenter in Zone J, and then followed by that in Zones H, D, C, and A, with the order described. This ordering reflects a progressive expansion of spatial flexibility, where each newly added datacenter provides additional siting options for AI companies to allocate their workloads.

### v3

@@ VI Case Study
We study an integrated system comprising the power grid, transportation network, and datacenter network. This case study illustrates how the framework developed in Section II can be applied to real-world settings, and examines how the resulting GCE captures the interactions and mutual influences among these networks and affects welfare outcomes. We evaluate the GCE in discrete one-hour market-clearing periods, treating all energy-use activities within each hour as simultaneous.
@@ VI-A Configurations of the Power and Multi-FL Systems
The power system configuration is adapted from the setting in [10636017], based on an 11-zone aggregation of the NYISO network that includes generation data, network parameters, baseline load profiles, and inter-zone distance information. The 11 zones, indexed by A-K, are shown in Fig. 4(a), where each zone represents a major operating region (e.g., WEST, CENTRL, N.Y.C), and the transmission lines connecting them indicate the network topology of the regional grid. The transportation and datacenter networks are electrically connected to the same power system, with configuration details illustrated below. To increase the sensitivity of LMPs to FL load-profile variations, we scale the quadratic coefficient in the generation cost by a factor of ten relative to [10636017], which steepens marginal generation costs.
@@ VI-A1 Electrified transportation system
The transportation network, shown in Fig. 4(b), is adapted from the Sioux Falls road network [SiouxFalls], consisting of 24 nodes and 76 edges. Each node is equipped with an EV charging station, while also representing the origins and destinations of EV trips. These nodes each connect electrically to one of the 11 NYISO zones in Fig. 4(a); the shaded blue areas indicate this geographic correspondence and determine $\mathbf{A}^{\mathrm{CB}}$. Arrowed lines indicate the bidirectional links between nodes, representing the road connections that form the transportation network.
@@
The EV-population data and OD-demand data are also adapted from [SiouxFalls], yielding 182 OD pairs with OD demand $\bm{\rho}$ and a total of 158,080 EV agents. To better capture real-world congestion effects, we rescale the OD demands. For each OD pair, we first identify the two routes with the lowest free-flow travel times. Given this fixed route set, charging is assigned to either the origin or the destination of each route, so that an EV’s decision trades off travel time against nodal charging prices. This construction yields a compact route–charging choice set that keeps the decision space low dimensional; allowing charging at intermediate locations would simply introduce additional routes without altering the modeling framework. The resulting route–charging configuration determines $\mathbf{A}^{\mathrm{CR}}$. The travel cost function on edge $e$ is modeled as $c_{e}(y_{e}(\mathbf{x}))=c_{e}^{0}\big(1+\kappa_{e}y_{e}(\mathbf{x})\big)$, where $c_{e}^{0}$ represents the base cost, computed as the free-flow travel time multiplied by a value of time of \$17.09/hour [van2025political]. The coefficient $\kappa_{e}$ reflects the marginal effect of traffic volume and is obtained by dividing the delay parameter by the edge capacity. Following [SiouxFalls], we triple the free-flow travel times and scale the delay parameters by a factor of ten to increase the sensitivity of per-driver travel cost to flow relocation. Each EV requires an identical charging amount $q^{\mathrm{EV}}\in[1,20]$ kWh, corresponding to the energy delivered during one hour of Level 1-2 charging [EVchargingspeed].
VI-A2 Datacenter network
The datacenter network consists of 11 AI companies and 5 datacenters. Figure 4(c) marks the geographic locations of all companies and datacenters, and links each of them to a power-zone index (A–K) in Fig. 4(a). We place the five datacenters in the WEST, CENTRL, NORTH, MILLWD, and N.Y.C zones following [10636017]; this placement defines the datacenter-to-bus coupling matrix $\mathbf{A}^{\mathrm{DB}}$. Each AI company can dispatch workload across all five datacenters; dashed lines in Fig. 4(c) illustrate this connectivity for one representative company, with different colors corresponding to different datacenters.
TABLE I: Parameters in AI Companies’ Utility Functions (23).
Param Description Value
$d_{j}^{(k)}$ Distance from company $k$ to datacenter $j$ based on [10636017]
$\zeta_{k}$ Total workload of user $k$ $10^{4}\cdot\mathcal{U}(1,5)$
The utility function of each AI company $k$ takes the quadratic form (23), where the parameters are specified as follows: Each component $\eta_{k,j}$ of the baseline utility vector represents company $k$’s willingness to pay (or preference) for datacenter $j$, denoted by $\tau^{(k)}_{j}$, minus a communication cost proportional to the distance $\gamma_{k}d^{(k)}_{j}$ between them. The quadratic cost term captures two effects. First, datacenter queueing [tsiligkaridis2025distributed] implies that the per-job waiting cost at datacenter $j$ increases with the aggregate workload $\sum_{\tilde{k}}x_{j}^{(\tilde{k})}$, scaled by a queueing coefficient $\alpha_{j}\geq 0$. As a result, company $k$ incurs a total waiting cost $\sum_{j}\alpha_{j}x_{j}^{(k)}\sum_{\tilde{k}}x_{j}^{(\tilde{k})}$, which couples the decisions of competing companies. Second, diminishing marginal returns for company $k$ are captured by an internal quadratic cost $\tfrac{1}{2}\omega_{k}\sum_{j}(x_{j}^{(k)})^{2}$ with $\omega_{k}>0$. These two effects yield the block structure of $\mathbf{M}$: for each $\tilde{k}\neq k$, the off-diagonal blocks satisfy $\mathbf{M}_{k,\tilde{k}}=\operatorname{diag}(\alpha_{1},\ldots,\alpha_{N^{\mathrm{D}}})$, and the row block $\mathbf{M}_{k,-k}$ concatenates $\{\mathbf{M}_{k,\tilde{k}}\}_{\tilde{k}\neq k}$. The diagonal blocks satisfy $\mathbf{M}_{k,k}=\operatorname{diag}(2\alpha_{1}+\omega_{k},\ldots,2\alpha_{N^{\mathrm{D}}}+\omega_{k})$, ensuring concavity of company $k$’s utility. These parameters of the utility functions are selected according to the aforementioned criteria and summarized in Table I, where $\mathcal{U}(\cdot)$ denotes the uniform distribution. Once the random components of the parameters in the table are determined, these parameters remain fixed throughout all experiments for consistency.
Each standardized workload (e.g., an AI inference/training batch) has an energy requirement of $q^{\mathrm{DC}}\in[0,10]$ kWh. Each AI company optimizes the allocation of its fixed total workload across datacenters to maximize its net payoff, and the total amount of workloads is 364,803.
@@ VI-B GCE-Based Interaction and Welfare Analysis
Using the GCE and SWM formulations developed in Section III, we compute the FL decisions, generator dispatch, and LMPs $(\mathbf{g}^{\star},\mathbf{s}^{\star},\bm{\lambda}^{\star})$ under GCE and compare them with the corresponding social welfare optimal outcomes. We then vary key parameters, namely, the transmission line capacities, the EV charging amount $q^{\mathrm{EV}}$, and the energy demand per workload $q^{\mathrm{DC}}$, to examine their impacts on the performance of each subsystem. For comparison, we consider two experimental settings: the congested case, based on the original dataset where limited line capacity causes congestion even without interconnection of the FL systems, and the uncongested case, a synthetic scenario in which all line-capacity constraints are relaxed to ensure an uncongested power system.
@@
We examine the impact of datacenter spatial flexibility by varying the number of datacenters that each AI company can access, from one to five, in Fig. 7, while keeping the total workload for each company fixed. Fewer available datacenters correspond to a more constrained workload-allocation decision space and less spatial flexibility. Concretely, we construct a nested sequence of access sets by progressively enabling connections from each company to additional datacenters: starting with the site in Zone J, then adding Zones H, D, C, and A. Thus, each step from 1 to 5 introduces one more siting option for workload placement, isolating the value of additional spatial flexibility. The one- and two-datacenter cases are infeasible because both sites lie within the power system’s most congested zone. In Fig. 7(a), increasing the number of datacenters decreases generation cost, as greater spatial flexibility enables workload redistribution toward more efficient and less congested regions of the grid. Spatial flexibility of the datacenter network likewise reduces EVs’ total travel cost under both GCE and SWM frameworks in Fig. 7(b) by flattening LMP disparities, thus mitigating traffic congestion in the transportation network induced by drivers chasing lower charging prices. As shown in Fig. 7(c), computing utility improves with spatial flexibility due to reduced queuing delays. When datacenter locations are restricted, workloads concentrate in a few sites, resulting in negative utility from excessive waiting. Overall, Fig. 7 demonstrates that expanding spatial flexibility benefits all coupled infrastructures even in a self-interested GCE model.
@@ VII Conclusion
Appendix A Non-Potential Game Case for Lemma 4
Theorem 3 (GCE for an Non-Potential Game).
[…]
Appendix B Proofs in Section III
@@
Since $J^{\mathrm{SW}}$ and $C$ are convex and $\mathcal{F}$ is closed and convex, the standard subgradient optimality condition for minimizing a convex function over a convex set (see [urruty1993convex]) implies $(\mathbf{s}^{\star},\mathbf{g}^{\star})$ is optimal for $\min_{(\mathbf{s},\mathbf{g})\in\mathcal{F}}J^{\mathrm{SW}}(\mathbf{s})+C(\mathbf{g})$ if and only if there exist $\bm{\mu}\in\partial J^{\mathrm{SW}}(\mathbf{s}^{\star}),\ \bm{\nu}\in\partial C(\mathbf{g}^{\star})$ such that
@@
Appendix C Proofs in Section IV
@@
which is exactly the UE condition (14). Conversely, any $\mathbf{x}^{\star}$ satisfying (14) admits such multipliers $\mu_{k},\nu_{r}$ and therefore satisfies the KKT conditions, proving necessity and sufficiency. This establishes the equivalence between the optimization problem (15) and the UE condition (14). A similar derivation appears in the classical transportation equilibrium theory [sheffi1985urban], where the charging component is absent. ∎
@@
TABLE II: Illustrative Example: Non-congested Case
@@
    No congestion $\bigl(f^{\max}\geq g_{2}^{\max}-(a-b)q/a\bigr)$. The line is not congested in both programs. Table II demonstrates the results under this parameter regime. We have $s_{1}^{\mathrm{GCE}}=2s_{1}^{\mathrm{SWM}}$ while the LMPs remain uniform, $\lambda_{1}^{\mathrm{GCE}}=\lambda_{2}^{\mathrm{GCE}}=\lambda_{1}^{\mathrm{SWM}}=\lambda_{2}^{\mathrm{SWM}}$. This result is an example for Lemma 3-i), where $(\mathrm{P}_{\mathrm{GCE}})$ does not support social welfare.
@@
Appendix D Proofs in Section V
@@
Since each utility $u_{k}$ is differentiable and concave and each feasible set $\mathcal{X}^{(k)}$ is convex and compact, the players’ optimization problem in (20) is convex. Its KKT conditions are therefore necessary and sufficient. Collecting the $K$ stationarity conditions yields exactly the VI in (21); Conversely, any $\mathbf{x}^{\star}$ solving (21) satisfies every player’s KKT system and is thus a Nash equilibrium. This one-to-one correspondence follows from Proposition 1.3.4 of [faechinei2003finite].
Now suppose the cross-partial symmetry condition (22) holds. By standard results in potential game theory [monderer1996potential], there exists a continuously differentiable scalar function $\Phi(\mathbf{x})$ satisfying $\nabla_{\mathbf{x}^{(k)}}\Phi=-\nabla_{\mathbf{x}^{(k)}}u_{k}$ for all $k$. Substituting this relation into (21) gives
@@
Appendix E Proof of Theorem 3
Proof of Theorem 3.
Fix an LMP vector $\bm{\lambda}\in\mathbb{R}^{N}$ and define the augmented operator $\mathbf{F}_{\bm{\lambda}}(\mathbf{x})\;:=\;\mathbf{F}(\mathbf{x})\;+\;\bigl(\mathbf{A}^{\mathrm{FL}}\bigr)^{\!\top}\bm{\lambda}.$ The VI problem (21) can be denoted as $\operatorname{VI}(\mathcal{X},\mathbf{F}_{\bm{\lambda}})$. Since $\mathbf{F}$ is strictly monotone on the compact, convex set $\mathcal{X}$, $\mathbf{F}_{\bm{\lambda}}$ is strictly monotone as well. Therefore, $\operatorname{VI}(\mathcal{X},\mathbf{F}_{\bm{\lambda}})$ has a unique solution according to Theorem 2.3.3 in [faechinei2003finite]. The merit function for $\operatorname{VI}(\mathcal{X},\mathbf{F}_{\bm{\lambda}})$ is as follows:
@@ Appendix E Proof of Theorem 3
Next, we illustrate the relationship between the characterization of GCE and the solution of $\operatorname{VI}(\Omega,\mathbf{F}_{\mathbf{C}})$. Since $C$ is convex and $\mathbf{F}$ is monotone, $\mathbf{F}_{\mathbf{C}}$ is a monotone operator on the convex set $\Omega$. By Proposition 1.3.4 of [faechinei2003finite], a pair $(\mathbf{x}^{\star},\mathbf{g}^{\star})$ solves $\operatorname{VI}(\Omega,\mathbf{F}_{\mathbf{C}})$ if and only if there exists a multiplier $\bm{\lambda}^{\star}$ associated with the coupling constraints $\mathbf{g}-\bm{\ell}-\mathbf{A}^{\mathrm{FL}}\mathbf{x}\!\in\!\mathcal{P}$ such that the triplet $(\mathbf{x}^{\star},\mathbf{g}^{\star},\bm{\lambda}^{\star})$ satisfies the KKT conditions for the VI, which consist of three parts: